OpenStudy (anonymous):

Salina earned $8 more than Tyler selling newspaper subscriptions. at the end of the day, Salina was given $1 as a sale bonus, and Tyler spent $3 on snacks. When they returned home, they had a total of $12. how much money did Tyler earn selling newspaper subscriptions?

OpenStudy (anonymous):

HELP!!!

OpenStudy (saifoo.khan):

call 911.

OpenStudy (anonymous):

gosh, comon i wanna finish ma hmmwk, its the last problem!!!

OpenStudy (anonymous):

6

OpenStudy (anonymous):

u sure about that?

OpenStudy (saifoo.khan):

8 + x +1 -3 = 12 x + 6 = 12 x = 12-6 x = 6

OpenStudy (anonymous):

yes am sure about it :)

OpenStudy (anonymous):

ohhh i see now..